Comparative Advantage and International Trade Portfolio

Millerville Jamestown Scooters Scooters 20 75 Helicopters Helicopters

The production possibilities curves above show all the possible combinations of helicopters and scooters that two towns, Millerville and Jamestown, can create using equal amounts of resources.

  1. Explain which country has the absolute advantage in the production of helicopters.

  2. Using the concept of opportunity cost that you learned in this lesson, explain which country has the comparative advantage in the production of scooters. You may include a table for your explanation. (Hint: Is this an input or output question?)

  3. If the two countries specialize and trade with each other, identify which country would import scooters. Explain.

  4. Explain if Jamestown would benefit from a trade with Millersville if the terms of trade were set such that one scooter is traded for one helicopter? Make sure to explain why they would benefit.
Answer #1

Ans)

Scooter Helicopter
Millerville 25 75
Jamestown 80 20

1) Absolute advantage is when one party can produce more goods than the other.

Here, Millerville has absolute advantage in Helicopter. While Jamestown has absolute advantage in scooter.

2) Comparative advantage is when one party can produce goods with lower opportunity cost.

For Millerville÷

Opportunity cost of scooter = helicopter sacrificed ÷ scooter gained = 75÷25 = 3

Opportunity cost of helicopter = scooter sacrificed ÷ helicopter gained = 25÷75 = 0.33

For Jamestown ÷

Opportunity cost of scooter = helicopter sacrificed ÷ scooter gained = 20÷80 = 0.25

Opportunity cost of helicopter = scooter sacrificed ÷ helicopter gained = 80÷20 = 4

Since Jamestown has lower opportunity cost of scooter, it has comparative advantage in scooter and will export scooter while Millerville will import scooter.

Since Millerville has less opportunity cost of helicopter, it has comparative advantage in helicopter and will export helicopter and Jamestown will import helicopters.

3) Countries trade if the exporting country gets more than the opportunity cost while importing country pays less than the opportunity cost.

Yes, they can benefit from trade because opportunity cost of scooter for Jamestown is less than 1 and opportunity cost of helicopter for Millerville is less than 1. That is, both the exporting parties are getting more than their opportunity cost.